OpenStudy (anonymous):

Find the limit. Lim approaches 0 from the right of (sin(x))/(8(sqrt(x)))

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{\sin x}{8 *\sqrt x}\]

OpenStudy (anonymous):

hmm maybe zero

OpenStudy (anonymous):

I think zero just glancing.

myininaya (myininaya):

satellite show isa your piecewise function

OpenStudy (anonymous):

\[f(x) = |x+4| = \left\{\begin{array}{rcc} x + 4 & \text{if} & x \geq -4 \\ - x - 4& \text{if} & x < -4 \end{array} \right. \]

OpenStudy (anonymous):

\[f(x)=\left\{ \begin{array}{rcc} x & \text{if}& x \ne 0\\ x^3 & \text{if}& x=0 \end{array} \right. \]